Changes between Version 48 and Version 49 of Status

Nov 30, 2009, 1:41:04 PM (14 years ago)



  • Status

    v48 v49  
    55Please contact Gary Llewellyn (Science / operations coordinator) if you have any questions or concerns: (01865 374 391)
     8== 30th November 2009 ==
     9''On Friday, LiDAR were collected from Abernethy forest (in support of GB09-11) data and today (Monday), at solar noon, LiDAR, photography and course resolution hyperspectral data were collected from Cumbria (in support of the current Urgency application, UR10-01). The collection of LiDAR data from Wales (in support of GB09-04) was also investigated but the area proved to be cloud covered and subject snow fall.''
     11== 24th November 2009 ==
     12''Last night the team were able to collect LiDAR data from the channel flood areas in Cumbria (UR10-01).  Minor cloud obscured some areas but the data were of a fine resolution (finer than the existing EA LiDAR datasets). Unfortunaely, cloud fully obscured the wider catchment. The aircraft landed at RAF Brize Norton (Gloucester and other local airports were closed) at ~00:30 hrs.   On return to Gloucster Airport these data will be downloaded, backed up and forwarded to our data processing team.''
     15== 23rd November 2009 ==
     16''Over the last week the ARSF have been waiting for clear conditions for several study sites. These were in London, Wales and Scotland. Support of GB08-19 (London) requires near solar noon clear condition, a preceding period of dry weather and clear, dry conditions for the following night. At this stage in the season GB09-04 (Coed-y-Brenin) and GB09-11 (Abernethy Forest) are basically LiDAR surveys (with photography if possible) and require clear condition between the ground and aircraft (night or day). These areas have not presented suitable opportunities. However, an additional data collection requirement was presented by the floods in Cumbria (UR10-01).  We launched on the weekend but were unsuccessful in our attempts to collect data from this site. A second attempt to collect data from the area is being made this evening in the hope of catching a potentially cloud free window between 21:00 and 00:00 hrs.''
    720== 26th October 2009 ==